### 11 Mar 2024 v1

@@ 4.2 Decoding and reproducing
It is possible to use inverse Fourier transform tools to retrieve the original message because the amplitude modulation behavior repeats itself periodically (infinite loop). In order to reproduce the message, either sounds (pressure waves) or laser pulses could be generated. One caveat of these approaches is fine-tuning the energy content of the reproduction. This is important to avoid under or over detection (the laser, for example, could be undetectable in a high energy - high frequency - scenario or could be as powerful as the solar system’s sun in a low energy - low frequency - scenario).
Another interesting approach is to use one-dimensional chains of trapped ions to produce a quantum mechanical system with discrete scale invariance and fractal-like time dependence [18]. In that case, it was shown that patterns similar to the Weierstrass function could be produced, which are pivotal to encode the messages using the framework hereby proposed. Considering the findings of Rose and Wright [19], sending such chains of trapped ions together with circuits to modulate the amplitude of the waves emitted could be an interesting transmission approach at hardware level. There are certainly limitations, but, in principle, for a certain range of amplitudes and frequencies, the transmission and emission could be achieved. Regarding software implementations, there are also limits up to which the waves can be generated and reproduced. These come from the limitations of precision for the calculations considered, for example.
@@
In the present paper, the messages being transmitted were binary streams representing, for example, ASCII encodings of known characters. By using fractal carriers, the message and the carriers became a single entity, which was shown to be a fractal. This suggests a more general framework where concepts could be stated in terms of fractals, which could ultimately encode not only characters, but also computation concepts in an integrated fractal way.
As a matter of fact, using fractals to establish a communication system may be beneficial as the universe has been shown to present fractal-like structures with remarkable self-similarity (the self-similar multifractal distributions of dark matter and gas, for example [20]). Therefore, the entities receiving the message could be familiar with self-similar structures, due to their occurrence in their own environment. Not only that, but literature indicates that digital sundials have been devised as projections of shadows of a fractal set [21], which would allow the possible communication of a correspondence between time scales related to Earth’s sun and the receivers’ sun. Therefore, using fractals as communication entities could be of interest.
Regarding possible relations between fractal geometry and computation, Penrose suggested that fractals might be a graphical way of looking at non-recursive mathematics [22]. Further exploring that rationale, Dube [23] established a relationship between the classical theory of computation and fractal geometry. In particular, he considered the concepts behind IFSs to define fractals, indicating that for every Turing machine (in the context of computation over integers and rationals) there exists a fractal set which can be viewed, in a certain sense, as geometrically encoding the complement of the language accepted by the machine. This suggests that in a fractal-based communication system, it would be possible to transmit fractal-based geometrical models of computation which are computationally universal. Inspired by previous works [21], Dube [23, 24] proposed a fractal-based “sun-computer”, indicating that for every Turing machine $M$ there exists such a fractal set $A$ which graphically “encodes” the behavior of $M$ on all inputs. This could be a potential carrier for a Turing Machine computational model without the need to use binary encodings.
@@ 5 Conclusion
We have introduced a new mathematical framework based on fractal theory that allows scale-independent communication.
The approach instantiates the idea of sending a self-encoding and self-decoding signal as a mathematical formula that carries both the data and its unfolding instructions. This would be similar to sending a self-executable message with its computer to run it. It is equivalent to sending self-executable computer code that, when unfolded, reads a message in all possible time scales and at the same time as if in all possible channels simultaneously.
Together with our previous work [5], we have contributed to specific technical challenges at the receiver and sender ends.

### 14 Mar 2024 v2

@@ 4.2 Decoding and reproducing
It is possible to use inverse Fourier transform tools to retrieve the original message because the amplitude modulation behaviour repeats itself periodically (infinite loop). In order to reproduce the message, either sounds (pressure waves) or laser pulses could be generated. One caveat when using these approaches is the need to fine-tune the energy content of the reproduction. This is important to avoid under- or over-detection (the laser, for example, could be undetectable in a high energy - high frequency scenario or could be as powerful as the solar system’s sun in a low energy - low frequency scenario).
Another interesting approach is to use one-dimensional chains of trapped ions to produce a quantum mechanical system with discrete scale invariance and fractal-like time dependence [18]. In such a case, it was shown that patterns similar to the Weierstrass function could be produced, which are pivotal to encode the messages using the framework proposed herein. In light of the findings of Rose and Wright [19], sending such chains of trapped ions together with circuits to modulate the amplitude of the waves emitted could be an interesting transmission approach at hardware level. There are certainly limitations, but in principle, for a certain range of amplitudes and frequencies, transmission and emission could be achieved. Regarding software implementations, there are also limits up to which the waves can be generated and reproduced. These arise from shortfalls in the precision of the calculations considered, for example.
@@
In the present paper, the messages being transmitted are binary streams representing, for example, ASCII encodings of known characters. Using fractal carriers, the message and the carriers became a single entity, which was shown to be a fractal. This suggests a more general framework where concepts could be stated in terms of fractals, which could ultimately encode not only characters but also computational concepts, in an integrated fractal way.
As a matter of fact, using fractals to establish a communication system may be beneficial as the universe has been shown to present fractal-like structures with remarkable self-similarity (the self-similar multifractal distributions of dark matter and gas, for example [20]). Therefore, the entities receiving the message could be familiar with self-similar structures, due to their occurrence in their own environment. Not only that, but the literature indicates that digital sundials have been devised as projections of shadows of a fractal set [21], which would allow the possible communication of a correspondence between time scales related to Earth’s sun and the receiver’s sun. Therefore, using fractals as communication entities could be of interest.
Regarding possible relations between fractal geometry and computation, Penrose has suggested that fractals might be a graphical way of looking at non-recursive mathematics [22]. Further exploring this rationale, Dube [23] established a relationship between the classical theory of computation and fractal geometry. In particular, he considered the concepts behind IFSs to define fractals, indicating that for every Turing machine (in the context of computation over integers and rationals) there exists a fractal set which can be viewed, in a certain sense, as geometrically encoding the complement of the language accepted by the machine. This suggests that in a fractal-based communication system, it would be possible to transmit fractal-based geometrical models of computation which are computationally universal. Inspired by previous works [21], Dube [23, 24] proposed a fractal-based “sun-computer”, indicating that for every Turing machine $M$ there exists such a fractal set $A$ which graphically “encodes” the behaviour of $M$ on all inputs. This could be a potential carrier for a Turing Machine computational model without the need to use binary encodings.
@@ 5 Conclusion
We have introduced a new mathematical framework based on fractal theory that allows spatio-temporal scale-independent and parallel infinite-frequency communication.
Our approach instantiates the idea of sending a self-encoding/self-decoding signal as a mathematical formula that carries both the data and its unfolding instructions. This would be similar to sending a self-executable message with a computer on which to run it. It is equivalent to sending self-executable computer code that, when unfolded, reads a message at all possible time scales and in all possible channels simultaneously.
One aspect to explore is how and whether we are prepared to receive and detect a signal with these properties. Together with our previous work [5], we have contributed to specific technical challenges at the receiver and now sender ends, and combining both approaches is to be done.